# MaNGA DynPop – VII. A Unified Bulge-Disk-Halo Model for Explaining Diversity in Circular Velocity Curves of 6000 Spiral and Early-Type Galaxies

### v1

@@ 2.2 Dynamical models, CVCs, and other dynamical properties
Figure 1: Examples of different shapes of JAM-derived CVCs (rising CVCs for the first row, flat CVCs for the second row, and declining CVCs for the third row). For each galaxy, the RGB image with plateifu (top left), the map of circular velocities (top middle), the map of line-of-sight velocity (top right), the CVC in the equatorial plane derived from JAM model (bottom left), the map of stellar age (bottom middle), the map of stellar metallicity (bottom right) are presented from left to right and top to bottom. The black contours are the observed surface brightness contours in steps of 1 mag. The black dots are the centroids of the Voronoi bins from which the maps were linearly interpolated.
Other structural and dynamical quantities used in this work are mainly taken from Zhu et al. (2023b), we briefly introduce the parameters and the corresponding keywords here:
@@
    where $F_{k}$, $V_{k}$, and $\sigma_{k}$ are the flux, stellar velocity, and stellar velocity dispersion in the $k$-th IFU spaxel. This quantity closely approximates the velocity dispersion $\sigma$ that one would measure by co-adding all spectra within the same aperture, and fitting $(V,\sigma)$ for that spectrum (Cappellari et al., 2013a, section 4.3).
@@
    where $F_{k}$, $V_{k}$ and $\sigma_{k}$ are the same as Equation (missing) 2; $R_{k}$ is the distance of $k$-th spaxel to the galaxy centre. The $\lambda_{\rm R_{e}}$ in the catalog of Zhu et al. (2023b) has been corrected for the beam smearing effect following Graham et al. (2018) [^2] https://github.com/marktgraham/lambdaR_e_calc.
@@
  • The ratio between scale radius of NFW halo and luminous half-light radius $r_{\rm s}/R_{\rm e}$: $r_{\rm s}$ is estimated from the Chabrier IMF-based stellar mass, which is converted from the Salpeter IMF-based stellar mass $M_{\rm\ast}$ (see the definition in Section 2.3) by subtracting 0.215 dex (Madau & Dickinson, 2014, figure 4), the stellar mass-to-halo mass relation in Moster et al. (2013), and the mass-concentration relation in Dutton & Macciò (2014).
@@ 2.3 Stellar population properties based on Stellar Population Synthesis (SPS)
    where $w_{i}$ is the fraction of mass contributed by the $i$-th template (the ppxf fit weight, if the templates are normalized to unitary mass), $L_{i}$ is the SDSS r-band luminosity per unit mass of the $i$-th template, and $x_{i}$ is the $\rm\lg Age$ (or $\mathrm{[Z/H]}$) of the $i$-th template. The radial gradients are measured by linearly fitting the $\langle\lg\rm Age\rangle$ or $\langle[Z/H]\rangle$ profiles within the elliptical half-light isophote.
@@
    where $M_{i}^{\rm nogas}$ is the stellar mass of the $i$-th template, which includes the mass of living stars and stellar remnants but excludes the mass of lost gas during stellar evolution.
@@
    where $(M_{\ast}/L)_{\rm SPS}$ is the SDSS r-band stellar mass-to-light ratio derived from the stacked spectrum within the elliptical half-light isophote (keyword: ML_int_Re), $L$ is the total luminosity derived from MGE model (keyword: Lum_tot_MGE).
@@ 3.1 The amplitude of CVCs
The amplitude of CVCs is usually characterized by the maximum (asymptotic) circular velocity (Tully & Fisher, 1977) for spirals, while the characteristic velocity measurement for ETGs is typically the velocity dispersion which enters the Faber-Jackson relation (Faber & Jackson, 1976) and Fundamental Plane (Djorgovski & Davis, 1987; Dressler et al., 1987). Previous studies aimed at unifying the dynamical relations of spirals and ETGs were done by converting the velocity dispersion $\sigma_{\rm e}$ into the circular velocity at half-light radius $V_{\rm circ}(R_{\rm e}^{\rm maj})$ or the maximum circular velocity $V_{\rm circ}^{\rm max}$ (Padmanabhan et al., 2004; Courteau et al., 2007; Schulz et al., 2010; Dutton et al., 2011; Cappellari et al., 2013a). Here, we revisit the empirical relations (in the form $V_{\rm circ}=k\sigma_{\rm e}$, where $k$ is the conversion factor) using the accurate $\sigma_{\rm e}$ provided by MaNGA (Law et al., 2021), together with the circular velocities derived from our dynamical models.
@@
The correlation between $\sigma_{\rm e}$ and $V_{\rm circ}(R_{\rm e}^{\rm maj})$ in MaNGA is presented in the top left panel of Figure 2. The best-fitting relation is given by
$\lg V_{\rm circ}(R_{\rm e}^{\rm maj})=2.36+0.87\times(\lg\sigma_{\rm e}-2.11),$ (8)
or approximately equals to
$V_{\rm circ}(R_{\rm e}^{\rm maj})\approx 3.31\times\sigma_{\rm e}^{0.87}\approx 1.65\times\sigma_{\rm e},$ (9)
@@
with an error of $\Delta/\sqrt{2}=9$%. When classifying the sample into different morphological types following Zhu et al. (2024, section 2.5), the conversion factors for ETGs ($k\approx 1.63$) and LTGs ($k\approx 1.71$) are similar. Comparisons with previous observation (ATLAS3D; Cappellari et al., 2013a), the EAGLE (Schaye et al., 2015; Crain et al., 2015) and Illustris-TNG (Naiman et al., 2018; Springel et al., 2018; Pillepich et al., 2018; Marinacci et al., 2018; Nelson et al., 2018) cosmological simulations (the relations taken from Ferrero et al., 2021), and theoretical prediction (Wolf et al., 2010) are shown in the top right panel. Cappellari et al. (2013a) found a slightly smaller factor $k\approx 1.51$ in ETGs, while Wolf et al. (2010) predicted $k=\sqrt{3}$. The relation in the EAGLE simulations also has a similar factor of $k\approx 1.72$ (or equivalently $V_{\rm circ}(R_{\rm e})=2.37\times\sigma_{\rm e}^{0.94}$). However, unlike the relations mentioned above, which show weak $\sigma_{\rm e}$ dependence, the relation in Illustris-TNG simulations ($V_{\rm circ}(R_{\rm e})=7.03\times\sigma_{\rm e}^{0.73}$) has an exponent significantly deviating away from one.
@@
A slightly tighter correlation between $\sigma_{\rm e}$ and $V_{\rm circ}^{\rm max}$ is observed in the bottom left panel of Figure 2, with
$\lg V_{\rm circ}^{\rm max}=2.37+0.93\times(\lg\sigma_{\rm e}-2.11),$ (10)
@@
$V_{\rm circ}^{\rm max}\approx 2.56\times\sigma_{\rm e}^{0.93}\approx 1.75\times\sigma_{\rm e},$ (11)
with an error of $\Delta/\sqrt{2}=8$%. No difference in the conversion factors are found for ETGs and LTGs (see the bottom left panel of Figure 2). This result is consistent with the conversion factor found for ETGs in ATLAS3D ($k\approx 1.76$; Cappellari et al., 2013a). Note that $V_{\rm circ}^{\rm max}$ is defined as the maximum circular velocity within the region where we have stellar kinematic data (usually 1.5 $R_{\rm e}$ or 2.5 $R_{\rm e}$ in MaNGA). When we only consider the galaxies whose radius of maximum circular velocity lies in the kinematic data range, the conversion factor $k\approx 1.75$ remains nearly unchanged but with a smaller error of $\Delta/\sqrt{2}=6$% (bottom right panel of Figure 2).
@@ 3.2 The shape of CVCs
The shape of CVCs also correlates with kinematic properties such as stellar velocity dispersion $\sigma_{\rm e}$ and the proxy for stellar angular momentum $\lambda_{\rm R_{e}}$ (first two panels in the second row). As $\sigma_{\rm e}$ increases and $\lambda_{\rm R_{e}}$ decreases, trends transition from rising CVCs to flat, and finally to declining CVCs. These trends reflect differences in galaxy structure: $\sigma_{\rm e}$ traces the bulge fraction (Cappellari et al., 2013b), while $\lambda_{\rm R_{e}}$ values reflect the fraction of ordered motions, which is also related to the prominence of the disk component or bulge mass fraction.
The $T_{50}$ and $T_{90}$ values represent the formation times of stars (in cosmic time) that account for 50% and 90% of the galaxy’s present-day stellar mass, respectively. The shape of CVCs shows clear dependence on $T_{90}$: declining CVCs for earlier quenched galaxies (higher $T_{90}$) and rising CVCs for later quenched galaxies (lower $T_{90}$). According to the two-phase evolutionary scenario on the mass-size plane (as discussed in Zhu et al., 2024), galaxies increase their mass and size through gas accretion-induced in situ star formation and dry mergers (by accreting ex situ already-formed stars). Earlier-quenched galaxies have already formed large bulges through in situ star formation and continue morphological transformation through dry mergers, while later-quenched galaxies still undergo bulge growth and transformation into earlier types. In contrast, there is little or no dependence on $T_{50}$ because $T_{50}$ is less relevant for distinguishing between the two mechanisms.
@@
These quantities often relate to bulge and disk fractions, which are seen as the physical origins of the diverse CVC shapes. One might assume it is possible to predict CVC shape given bulge and disk fractions, and dark matter fraction that accounts for the dark matter halo contribution. In the next section, we will assess whether this model (with a bulge, a disk, and a dark matter halo) can reproduce the observed diverse CVC shapes.
Figure 3: The normalized CVCs correlate with other galaxy properties. Definitions can be found in Section 2.2 and Section 2.3. These properties include masses (dynamical mass $M_{\rm JAM}$, stellar mass $M_{\ast}$ based on Salpeter IMF), morphology (T-Type and $B/T$), velocity dispersion $\sigma_{\rm e}$, spin parameter $\lambda_{\rm R_{e}}$, the cosmic times when stars accounting for 50% ($T_{50}$) and 90% ($T_{90}$) of galaxies’ present-day stellar mass were formed, luminosity-weighted stellar age, stellar metallicity, and their radial gradients.
3.3 Bulge-disk-halo model to quantify the shape of CVCs
Based on the remarkable similarity of RCs first noted by Rubin et al. (1985), there have been some studies trying to predict the amplitude and shape of RCs with a few key photometric parameters. For instance, Persic et al. (1996) and Salucci et al. (2007) proposed a universal rotation curve (URC) model for spiral galaxies, and Karukes & Salucci (2017) and Di Paolo et al. (2019) extended the URC model to dwarf and low surface brightness galaxies, respectively. Recently, Patel et al. (2024) developed a neural network (NN) model to predict the shape of RCs using the same photometric quantities as the literature URC, and found higher accuracy with their NN-based URC. They concluded that the improvement of literature URC at all radii requires detailed modeling in the inner region (including a bulge component) and at the outskirts (better parameterization of dark matter halo). The inclusion of a bulge component is even more important when quantifying the shape of CVCs in MaNGA galaxies, which include a large fraction of bulge-dominated systems.
In this section, we try to use a bulge-disk-halo (hereafter BDH) model to predict the shape of CVCs. In this model, the bulge and disk components are described as two Sersic components, with the Sersic indices for the bulge and disk being $n_{\rm Ser,bulge}$ and $n_{\rm Ser,disk}=1$ (exponential), respectively, while the dark matter halo is assumed to be an NFW profile. For each Sersic component, the surface brightness is written as

### v2

@@ 2.2 Dynamical models, CVCs, and other dynamical properties
Table 2: Systematic errors of circular velocity measurements
Quantities Qual = 0 Qual = 1 Qual = 2 Qual = 3 Mass Models
Slope $\Delta$ Error Slope $\Delta$ Error Slope $\Delta$ Error Slope $\Delta$ Error
(dex) (dex) (dex) (dex)
$V_{\rm circ}(R_{\rm e})$ 1.00 0.022 3.65% 1.00 0.0098 1.61% 1.00 0.018 2.97% 1.00 0.0047 0.77% NFW, gNFW
$V_{\rm circ}(R_{\rm e}^{\rm maj})$ 1.00 0.020 3.31% 1.00 0.0099 1.62% 1.01 0.018 2.97% 1.00 0.0060 0.98% NFW, gNFW
$V_{\rm circ}^{\rm max}$ 1.00 0.022 3.65% 1.00 0.011 1.81% 1.01 0.018 2.97% 1.00 0.0056 0.92% NFW, gNFW
$V_{\rm circ}({\rm rmax})$ 1.01 0.024 3.98% 1.00 0.011 1.81% 1.01 0.018 2.97% 1.00 0.0061 1.00% NFW, gNFW
Note. — The systematic errors of circular velocities ($V_{\rm circ}(R_{\rm e})$, $V_{\rm circ}(R_{\rm e}^{\rm maj})$, $V_{\rm circ}^{\rm max}$, $V_{\rm circ}({\rm rmax})$) for different quality groups, following the calculation presented in table 3 of Zhu et al. (2023b). The slope and $\Delta$ denote the slope and observed scatter obtained by the lts_linefit procedure (with clip=6). The error is defined as $\Delta/\sqrt{2}$ assuming that the quantities on both axes are comparable. The errors of quantities in this table are derived from four models (i.e. the more flexible NFW and gNFW models, each with two orientations of the velocity ellipsoid).
Figure 1: Examples of different shapes of JAM-derived CVCs (rising CVCs for the first two rows, flat CVCs for the second two rows, and declining CVCs for the third two rows). For each galaxy, there are six images: the RGB image with plateifu (top left), the map of circular velocities (top middle), the map of line-of-sight velocity (top right), the CVC in the equatorial plane derived from JAM model (bottom left), the map of stellar age (bottom middle), the map of stellar metallicity (bottom right). The black contours are the observed surface brightness contours in steps of 1 mag. The black dots are the centroids of the Voronoi bins from which the maps were linearly interpolated.
Other structural and dynamical quantities used in this work are mainly taken from Zhu et al. (2023b), which are derived directly from observational data and the best-fitting dynamical models (with five free parameters for the most flexible gNFW model). We briefly introduce the parameters and the corresponding keywords here:
@@
    where $F_{k}$, $V_{k}$, and $\sigma_{k}$ are the flux, stellar velocity, and stellar velocity dispersion in the $k$th IFU spaxel. This quantity closely approximates the velocity dispersion $\sigma$ that one would measure by coadding all spectra within the same aperture, and fitting $(V,\sigma)$ for that spectrum (Cappellari et al., 2013a, section 4.3).
@@
    where $F_{k}$, $V_{k}$ and $\sigma_{k}$ are the same as Equation () 2; $R_{k}$ is the distance of $k$th spaxel to the galaxy center. The $\lambda_{\rm R_{e}}$ in the catalog of Zhu et al. (2023b) has been corrected for the beam-smearing effect following Graham et al. (2018) [^2] https://github.com/marktgraham/lambdaR_e_calc.
@@
  • The ratio between the scale radius of the NFW halo and the luminous half-light radius $r_{\rm s}/R_{\rm e}$: $r_{\rm s}$ is estimated from the Chabrier IMF-based stellar mass, which is converted from the Salpeter IMF-based stellar mass $M_{\rm\ast}$ (see the definition in Section 2.3) by subtracting 0.215 dex (Madau & Dickinson, 2014, figure 4), the stellar mass-to-halo mass relation in Moster et al. (2013), and the mass-concentration relation in Dutton & Macciò (2014).
@@ 2.3 Stellar population properties based on Stellar Population Synthesis (SPS)
    where $w_{i}$ is the fraction of mass contributed by the $i$th template (the ppxf fit weight, if the templates are normalized to unitary mass), $L_{i}$ is the SDSS r-band luminosity per unit mass of the $i$th template, and $x_{i}$ is the $\rm\lg Age$ (or $\mathrm{[Z/H]}$) of the $i$th template. The radial gradients are measured by linearly fitting the $\langle\lg\rm Age\rangle$ or $\langle[Z/H]\rangle$ profiles within the elliptical half-light isophote.
@@
    where $M_{i}^{\rm nogas}$ is the stellar mass of the $i$th template, which includes the mass of living stars and stellar remnants but excludes the mass of lost gas during stellar evolution.
@@
    where $(M_{\ast}/L)_{\rm SPS}$ is the SDSS $r$-band stellar mass-to-light ratio derived from the stacked spectrum within the elliptical half-light isophote (keyword: ML_int_Re), and $L$ is the total luminosity derived from MGE model (keyword: Lum_tot_MGE).
@@ 3.1 The amplitude of CVCs
The amplitude of CVCs is usually characterized by the maximum (asymptotic) circular velocity (Tully & Fisher, 1977) for spirals, while the characteristic velocity measurement for ETGs is typically the velocity dispersion which enters the Faber-Jackson relation (Faber & Jackson, 1976) and fundamental plane (Djorgovski & Davis, 1987; Dressler et al., 1987). Previous studies aimed at unifying the dynamical relations of spirals and ETGs were done by converting the velocity dispersion $\sigma_{\rm e}$ into the circular velocity at half-light radius $V_{\rm circ}(R_{\rm e}^{\rm maj})$ or the maximum circular velocity $V_{\rm circ}^{\rm max}$ (Padmanabhan et al., 2004; Courteau et al., 2007; Schulz et al., 2010; Dutton et al., 2011; Cappellari et al., 2013a). Here, we revisit the empirical relations (in the form $V_{\rm circ}=k\sigma_{\rm e}$, where $k$ is the conversion factor) using the accurate $\sigma_{\rm e}$ provided by MaNGA (Law et al., 2021), together with the circular velocities derived from our dynamical models. To avoid potential overestimation of circular velocities (or total masses) for nearly face-on galaxies (e.g. fig. 12 in Lablanche et al., 2012), we further exclude those galaxies with inclination angles below 30 (accounting for $\sim 15\%$ of galaxies with Qual $\geqslant 1$) in this section. This exclusion has a negligible impact on the best-fitting relations, altering the coefficients and scatter by only a few percent.
@@
The correlation between $\sigma_{\rm e}$ and $V_{\rm circ}(R_{\rm e}^{\rm maj})$ in MaNGA is presented in the top-left panel of Figure 2. The best-fitting relation is given by
$\lg V_{\rm circ}(R_{\rm e}^{\rm maj})=2.35+0.86\times(\lg\sigma_{\rm e}-2.11),$ (8)
or approximately equal to
$V_{\rm circ}(R_{\rm e}^{\rm maj})\approx 3.43\times\sigma_{\rm e}^{0.86}\approx 1.62\times\sigma_{\rm e},$ (9)
@@
with an error of $\Delta/\sqrt{2}=7$%. When classifying the sample into different morphological types following Zhu et al. (2024, section 2.5), the conversion factors for ETGs ($k\approx 1.59$) and LTGs ($k\approx 1.71$) are similar. Comparisons with previous observation (ATLAS3D; Cappellari et al., 2013a), the EAGLE (Schaye et al., 2015; Crain et al., 2015) and Illustris-TNG (Naiman et al., 2018; Springel et al., 2018; Pillepich et al., 2018; Marinacci et al., 2018; Nelson et al., 2018) cosmological simulations (the relations taken from Ferrero et al., 2021), and theoretical prediction (Wolf et al., 2010) are shown in the top-right panel. Cappellari et al. (2013a) found a slightly smaller factor $k\approx 1.51$ in ETGs, while Wolf et al. (2010) predicted $k=\sqrt{3}$. The relation in the EAGLE simulations also has a similar factor of $k\approx 1.72$ (or equivalently $V_{\rm circ}(R_{\rm e})=2.37\times\sigma_{\rm e}^{0.94}$). However, unlike the relations mentioned above, which show weak $\sigma_{\rm e}$ dependence, the relation in Illustris-TNG simulations ($V_{\rm circ}(R_{\rm e})=7.03\times\sigma_{\rm e}^{0.73}$) has an exponent significantly deviating away from 1.
@@
A slightly tighter correlation between $\sigma_{\rm e}$ and $V_{\rm circ}^{\rm max}$ is observed in the bottom-left panel of Figure 2, with
$\lg V_{\rm circ}^{\rm max}=2.36+0.92\times(\lg\sigma_{\rm e}-2.11),$ (10)
@@
$V_{\rm circ}^{\rm max}\approx 2.62\times\sigma_{\rm e}^{0.92}\approx 1.72\times\sigma_{\rm e},$ (11)
with an error of $\Delta/\sqrt{2}=7$%. No significant difference in the conversion factors are found for ETGs ($k\approx 1.71$) and LTGs ($k\approx 1.75$), as shown in the bottom-left panel of Figure 2. This result is consistent with the conversion factor found for ETGs in ATLAS3D ($k\approx 1.76$; Cappellari et al., 2013a). Note that $V_{\rm circ}^{\rm max}$ is defined as the maximum circular velocity within the region where we have stellar kinematic data (usually 1.5 $R_{\rm e}$ or 2.5 $R_{\rm e}$ in MaNGA). When we only consider the galaxies whose radius of maximum circular velocity lies in the kinematic data range, the conversion factor $k\approx 1.72$ remains nearly unchanged but with a smaller error of $\Delta/\sqrt{2}=5$% (bottom-right panel of Figure 2).
@@ 3.2 The shape of CVCs
The shape of CVCs also correlates with kinematic properties such as stellar velocity dispersion $\sigma_{\rm e}$ and the proxy for stellar angular momentum $\lambda_{\rm R_{e}}$ (first two panels in the second row of Figure 3). As $\sigma_{\rm e}$ increases and $\lambda_{\rm R_{e}}$ decreases, trends transition from rising CVCs to flat, and finally to declining CVCs. These trends reflect differences in galaxy structure: $\sigma_{\rm e}$ traces the bulge fraction (Cappellari et al., 2013b), while $\lambda_{\rm R_{e}}$ values reflect the fraction of ordered motions, which is also related to the prominence of the disk component or bulge mass fraction.
The $T_{50}$ and $T_{90}$ values represent the formation times of stars (in cosmic time) that account for 50% and 90% of the galaxy’s present-day stellar mass, respectively. The shape of CVCs shows clear dependence on $T_{90}$: declining CVCs for earlier-quenched galaxies (higher $T_{90}$) and rising CVCs for later-quenched galaxies (lower $T_{90}$). According to the two-phase evolutionary scenario on the mass-size plane (as discussed in Zhu et al., 2024), galaxies increase their mass and size through gas accretion-induced in situ star formation and dry mergers (by accreting ex situ already formed stars). Earlier-quenched galaxies have already formed large bulges through in situ star formation and continue morphological transformation through dry mergers, while later-quenched galaxies still undergo bulge growth and transformation into earlier types. In contrast, there is little or no dependence on $T_{50}$ because $T_{50}$ is less relevant for distinguishing between the two mechanisms.
@@
These quantities often relate to bulge and disk fractions, which are seen as the physical origins of the diverse CVC shapes. One might assume it is possible to predict CVC shape given bulge and disk fractions, and a DM fraction that accounts for the DM halo contribution. In the next subsection, we will assess whether this model (with a bulge, a disk, and a DM halo) can reproduce the observed diverse CVC shapes.
Figure 3: The normalized CVCs correlate with other galaxy properties. Definitions can be found in Section 2.2 and Section 2.3. These properties include masses (dynamical mass $M_{\rm JAM}$ and stellar mass $M_{\ast}$ based on Salpeter IMF), morphology (T-Type and $B/T$), velocity dispersion $\sigma_{\rm e}$, spin parameter $\lambda_{\rm R_{e}}$, the cosmic times when stars accounting for 50% ($T_{50}$) and 90% ($T_{90}$) of galaxies’ present-day stellar mass were formed, luminosity-weighted stellar age, stellar metallicity, and their radial gradients.
3.3 A Bulge-Disk-Halo model to quantify the shape of CVCs
Based on the remarkable similarity of RCs first noted by Rubin et al. (1985), there have been some studies trying to predict the amplitude and shape of RCs with a few key photometric parameters. For instance, Persic et al. (1996) and Salucci et al. (2007) proposed a universal rotation curve (URC) model for spiral galaxies, and Karukes & Salucci (2017) and Di Paolo et al. (2019) extended the URC model to dwarf and low-surface-brightness galaxies, respectively. Recently, Patel et al. (2024) developed a neural network (NN) model to predict the shape of RCs using the same photometric quantities as the literature URC, and found higher accuracy with their NN-based URC. They concluded that the improvement of literature URC at all radii requires detailed modeling in the inner region (including a bulge component) and at the outskirts (better parameterization of a DM halo). The inclusion of a bulge component is even more important when quantifying the shape of CVCs in MaNGA galaxies, given their substantial population of bulge-dominated systems that has been demonstrated through combined photometric and kinematic structural decompositions (Rigamonti et al., 2023, 2024).
In this section, we try to use a bulge-disk-halo (hereafter BDH) model to predict the shape of CVCs. In this model, the bulge and disk components are described as two Sersic components, with the Sersic indices for the bulge and disk being $n_{\rm Ser,bulge}$ and $n_{\rm Ser,disk}=1$ (exponential), respectively, while the DM halo is assumed to be an NFW profile. For each Sersic component, the surface brightness is written as